Anya and Ivan lean over a balcony side by side. Anya throws a penny downward with an initial speed of 5 m/s. Ivan throws a penny upward with the same speed. Both pennies end up on the ground below. Compare their kinetic energies and velocities on impact.


1
Expert's answer
2021-09-01T11:20:30-0400

The speed of the coin thrown by Anya and Ivan:


"v_I=\sqrt{v_0-2gh},\\
v_A=\sqrt{2gh+v_0.}"


As we see, the speed of Anya's coin will be greater.

Find their kinetic energies:


"E_I=\frac12 mv_I^2=\frac12m(v_0-2gh),\\\space\\
E_A=\frac12mv_A^2=\frac12m(v_0+2gh)."

The kinetic energy of Anya's coin is greater as well .